OpenStudy (anonymous):

Find all real zeros of y=x^4 - 40x^2 + 144

OpenStudy (cwrw238):

let x^2 = X solve X^3 - 40X + 144 = 0

OpenStudy (cwrw238):

then 2 values of X give x = +/- sqrtX

OpenStudy (cwrw238):

* thats X^2 not X^3

OpenStudy (cwrw238):

you can solve X^2 - 40X + 144 = 0 by factoring

OpenStudy (cwrw238):

can you solve that equation? do you need help?

OpenStudy (anonymous):

I came up with x=+-2 and x=+-6 are those correct?

OpenStudy (cwrw238):

yes